OpenStudy (amurciano2):

Can someone help me with this calc2 questions integral of x5^x dx. I know that you have to integrate by parts, but I am not sure how do to that with this problem.

OpenStudy (irishboy123):

you'll need this \((5^x)' = 5^x .ln5 \) or \(\int 5^x \ dx = \ \frac{1}{ln5}.5^x\) so set it up with \(u = x \) and \( v' = 5^x\) and plough away ask if help needed :p

OpenStudy (dinamix):

\[x *\exp (x \ln5) \] will be easy for integral by parts @IrishBoy123

OpenStudy (irishboy123):

sweet as a nut, @dinamix !!

OpenStudy (dinamix):

i gave him easy metohd Mr. irish

OpenStudy (dinamix):

lol

OpenStudy (amurciano2):

How do you integrate 5^x/ln5

OpenStudy (irishboy123):

\(\large \int \frac{5^x}{ln5} = \frac {1}{ln5}\int 5^x\) and i posted that integral in my first post :p